OpenStudy (anonymous):

use the unit ciecle find value sec(-90°)

OpenStudy (anonymous):

\[\sec(-90) = \frac{1}{\cos(-90)}= \frac{1}{0}\] so it is undefined.

OpenStudy (anonymous):

ty